# COOL: Efficient and Reliable Chain-Oriented Objective Logic with Neural Networks Feedback Control

### v2

@@ Abstract
Program synthesis methods, whether formal or neural-based, lack fine-grained control and flexible modularity, which limits their adaptation to complex software development. These limitations stem from rigid Domain-Specific Language (DSL) frameworks and neural network incorrect predictions. To this end, we propose the Chain of Logic (CoL), which organizes synthesis stages into a chain and provides precise heuristic control to guide the synthesis process. Furthermore, by integrating neural networks with libraries and introducing a Neural Network Feedback Control (NNFC) mechanism, our approach modularizes synthesis and mitigates the impact of neural network mispredictions. Experiments on relational and symbolic synthesis tasks show that CoL significantly enhances the efficiency and reliability of DSL program synthesis across multiple metrics. Specifically, CoL improves accuracy by 70% while reducing tree operations by 91% and time by 95%. Additionally, NNFC further boosts accuracy by 6%, with a 64% reduction in tree operations under challenging conditions such as insufficient training data, increased difficulty, and multidomain synthesis. These improvements confirm COOL as a highly efficient and reliable program synthesis framework.
Figure 1: Chain-of-Logic (highlighted part) organizes the rule application into a structured sequence, enhancing the Domain-Specific Language (DSL) framework’s ability to handle complex tasks. The Neural Network Feedback Control mechanism (red path) utilizes data during synthesis to improve the performance of the synthesis process dynamically.
@@ 1 Introduction
In this paper, following the principles of fine-grained control and flexible modularity, we present COOL (Chain-Oriented Objective Logic), a neural-symbolic framework for complex program synthesis. At the core of our approach, we introduce the Chain-of-Logic (CoL), which integrates the functions of the activity diagram to enable fine-grained control Gomaa (2011). As illustrated in Figure 1, programmers can precisely organize rules into multiple stages and manage control flow using heuristics and keywords. Additionally, we leverage neural networks on top of CoL to dynamically fine-tune the synthesis process. For this purpose, we introduce Neural Network Feedback Control (NNFC) Turan & Jäschke (2024), which enhances future synthesis by learning from data generated during synthesis and suppresses neural network incorrect predictions through filtering. To ensure modularity, each neural network is bound with a specific CoL DSL, stored in separate library files for clear isolation and easy reuse. Thus, through the combination of CoL and NNFC, COOL achieves high efficiency and reliability when tackling complex synthesis tasks.
@@
Figure 3: Chain-of-Logic. In this illustrative CoL DSL, each node represents a stage or activity where a set of rules can be applied to generate partial programs. The flow between stages is managed by keywords return, logicjump(n), and abort, allowing for the implementation of complex control flow in program synthesis.
@@ 2.1 Chain-of-Logic (CoL)
Activity diagrams, widely used in software engineering, effectively describe how an initial state transitions to a final state through multiple stages. This feature aligns with the DSL-based program synthesis process. A DSL, defined as a context-free grammar, converts partial programs with nonterminal symbols into complete programs by applying given rules. However, as the rule set grows, DSL becomes inefficient in exploring partial programs. To enhance the efficiency of DSL, the Chain-of-Logic, drawing inspiration from activity diagrams, organizes rule applications during synthesis into a sequence of manageable stages, as illustrated in 3.
CoL improves the control flow of the DSL with two key features: heuristic vectors and keywords. Heuristic vectors specify the stages where rules apply and their corresponding values. For example, in Figure 1, a rule with the heuristic vector (0,7,3) is applicable in stages 2 and 3 with heuristic values of 7 and 3, respectively. These vectors form the core of CoL’s control flow.
Second, CoL introduces three keywords—return, logicjump(n), and abort—to dynamically choose the next stage during synthesis:
@@
    return: Ends the current rule, staying within current stage or advancing to following stages.
@@
    logicjump(n): Jumps directly to the stage $n$, enabling branching and loops within CoL.
@@
In summary, CoL provides fine-grained control through heuristic vectors and keywords. This structured and detailed approach enhances the efficiency of DSL synthesis.
Figure 4: Neural Network Feedback Control. The left side illustrates the complete control loop of NNFC. In the forward flow (green path), heuristic values $u$ guide the synthesis process as control signals. In the feedback loop (red path), the DSNN (Domain-Specific Neural Network, the neural network paired with a DSL) generates initial error signals $e_{0}$ from partial programs $y$. These singals are then filtered to produce high-quality error signals $e_{1}$, which adjust the initial heuristic values $u_{0}$. In multidomain synthesis, the CoL DSL and DSNN from the self-domain use partner domain information (dashed path) to clarify tasks and avoid competition, ensuring modularity. The right side details the feedback loop: The DSNN comprises multiple neural networks coupled in series via noise signals, with each network generating its own error signal $e_{0}$, then these signals with large discrepancies are filtered, retaining the final high-quality error signals $e_{1}$.
@@ 2.2 Neural Network Feedback Control (NNFC)
The quality of the signals generated in the feedback loop directly determines the effectiveness of NNFC. If the error signals are of poor quality, NNFC may not only fail to provide additional improvements but also degrade CoL DSL performance. We ensure the error signal quality through an inner coupling structure within DSNN. As shown in Figure 4 (right), during synthesis tasks, DSNN processes partial programs using a series of sequentially connected neural networks. Each neural network takes both the partial programs and intermediate results from the preceding neural network as input, generating its own predictions. When errors occur in earlier networks, they propagate downstream as noise signals, amplifying at each stage. The difference in the outputs between these neural networks is positively correlated with the accumulated error. To mitigate this, we set a threshold to filter out signals with a significant difference in outputs. Finally, DSNN uses passed signals to generate multi-head outputs to fine-tune the forward flow:
@@ 3.1 Experimental Setup
Benchmarks. We evaluate CoL and NNFC using relational and symbolic tasks with varying difficulty levels, as detailed in Table 1. Specifically, the relational tasks are drawn from the CLUTRR Sinha et al. (2019) dataset, where the goal is to synthesize programs that capture specific target relationships based on human common-sense reasoning. In contrast, the symbolic tasks are generated by GPT Achiam et al. (2023). They involve synthesizing standard quadratic equation programs from non-standard quadratic forms by performing manual calculation steps. Although these tasks are simple for humans, they serve as a straightforward demonstration of how fine-grained control, derived from programmer expertise, can significantly improve program synthesis efficiency.
Metrics. Besides accuracy, we also focus on the following points: (1) CPU Overhead is assessed by the number of tree operations required for synthesis. (2) Memory overhead is assessed by the number of transformation pairs (a partial program paired with the rule to be applied) [^1] Each partial program must be completed with at most 1000 transformation pairs, though this may exceed 1000 if additional tasks are generated during synthesis.. (3) GPU Overhead is measured by the number of neural network invocations. (4) Time overhead is referenced by the actual time spent on program synthesis tasks. (5) Filtering Performance is evaluated by the attenuation ratio of invalid to passed neural network predictions.
@@
Table 2: CoL DSL configurations. The DSL for relational benchmarks has a limited search space and shorter CoL, facing challenges from numerous production rules leading to larger trees. Conversely, the DSL for symbolic benchmarks offers an unlimited search space with a longer CoL, but the many permutation rules increase the risk of cyclic rule applications.
@@ 3.3 Dynamic Experiments
Further ablation experiments confirm that reliability provided by NNFC primarily stems from the filtering effect of the inner coupling structure. As shown in Figures 6 and 7, the inner coupling structure reduces the occurrence of accuracy declines due to DSNN mispredictions by 94%. Additionally, the dynamic performance reveals how the inner coupling structure enhances NNFC:
In the scenarios where a DSNN underperforms due to issues such as insufficient training data Mikołajczyk & Grochowski (2018) (as seen in Figure 6, tasks 51-100), inadequate generalization to more challenging tasks Yosinski et al. (2014); Wei et al. (2019) (Figure 6, tasks 301-350), and catastrophic forgetting when tasks from a new domain are learned Kirkpatrick et al. (2017); Van de Ven & Tolias (2019) (Figure 7, tasks 1-100), incorrect predictions lead the actual synthesis path to deviate from the CoL, which in turn causes inefficiency and reduced accuracy. During these phases, for NNFC with the inner coupling structure, the attenuation ratio spikes, indicating that a large percentage of neural network predictions are filtered out. Consequently, the inner coupling structure ensures that the synthesis process adheres to the Course of Learning (CoL), effectively mitigating the negative impact of DSNN mispredictions and enhancing reliability.
@@ 5 Conclusion
Appendix A Rule in CoL DSL
Appendix B Stage Progression Driven by Heuristic Vectors
Let $s$ denote the CoL stage, $h$ donate the heuristic value, and $n$ donate the length of CoL. A rule’s heuristic vector can be mathematically represented as:
$\mathbf{H}=$$\{(s_{0},h_{0}),(s_{1},h_{1}),\ldots,(s_{n},h_{n})\},\quad n\in\mathbb{N}^{+}$ (1)
Upon applying a rule with heuristic vector $\mathbf{H}$, the subsequent stage, $s_{\text{next}}$, can only advance or remain the same, and the next stage should be as close to the current stage as possible:
$\min{s_{\text{next}}}\quad\text{such that}\quad\exists$ $(s_{\text{next}},h_{\text{next}})\in\mathbf{H}\quad\text{and}\quad s_{\text{next}}\geq s_{\text{current}}$ (2)
Appendix C Neural Networks in DSNN
COOL performs synthesis tasks using Three-Address Code (TAC), also utilized as input by DSNN. TAC serves as an intermediate representation (IR), allowing program synthesis to be conducted without the constraints of specific DSL syntax or the machine code format of the execution platform Sujeeth et al. (2014). As TAC embodies both the graphical properties of a syntax tree and the sequential properties of execution, the design of the neural network must be capable of capturing these dual characteristics.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Figure 4 (right) illustrates using three neural network units arranged in series to construct the internal coupling structure of DSNN. Labeling these neural networks with A, B, and C in order of their sequence, Table 6 details the specific input features for each network: Neural network B receives its input feature ”applied” from network A’s output feature ”jumps,” while network C’s input features ”applied” and ”next stage” are derived from the output features ”jumps” and ”next stage” of network B. The output features of three neural network units are consistent and comparable, Table 7 presents the output features of the neural networks.
It is necessary to note that the DSNNs without internal coupling structures in Table 3 contain only neural network A.
Feature Feature Size Neural Network Signification
grounded 2 A, B, C The node is in a fully specified expression.
root 2 A, B, C The tree representing the subtask is rooted at this node.
non-terminal 2 A, B, C The node is a non-terminal.
type 1 A, B, C Type of the node.
identifier 1 A, B, C Identifier of the node.
string 1 A, B, C The node contains a string as the immediate value.
number 1 A, B, C The node contains a number as the immediate value.
operator 1 A, B, C The node is an operator.
applied 1 B, C A rule is applied to the subtree rooted at this node (derived from the output feature “jumps” of the previous neural network).
next stage 1 C The CoL stage to advance to after applying the rule (derived from the output feature “next stage” of the previous neural network).
Feature Feature Size Neural Network Signification
domain 2 A, B, C Relevance of task domains to DSNN.
feasibility 2 A, B, C Feasibility of synthesizing the complete program.
jumps max_tree_depth*3 A, B, C The path from the tree’s root to the subtree’s root where the rule is applied (jump left, right, or stop in each step).
next stage 1 A, B, C The CoL stage to advance to after applying the rule.
heuristic sign 2 A, B, C Sign of the rule’s heuristic value.
heuristic value 1 A, B, C Rule’s heuristic value.
expression 2 A, B, C Type of rule’s head (expression or terminal).
Appendix D Signal Clipper
D.1 A* Search in Program Synthesis

### v3

@@ Abstract
Program synthesis methods, whether formal or neural-based, lack fine-grained control and flexible modularity, which limits their adaptation to complex software development. These limitations stem from rigid Domain-Specific Language (DSL) frameworks and neural network incorrect predictions. To this end, we propose the Chain of Logic (CoL), which organizes the synthesis process into an activity flow and provides heuristic control to guide the process. Furthermore, by integrating neural networks with libraries and introducing a Neural Network Feedback Control (NNFC) mechanism, our approach modularizes synthesis and mitigates the impact of neural network mispredictions. Experiments on relational and symbolic synthesis tasks show that CoL significantly enhances the efficiency and reliability of DSL program synthesis across multiple metrics. Specifically, CoL improves accuracy by 70% while reducing tree operations by 91% and time by 95%. Additionally, NNFC further boosts accuracy by 6%, with a 64% reduction in tree operations under challenging conditions such as insufficient training data, increased difficulty, and multidomain synthesis. These improvements confirm COOL as a highly efficient and reliable program synthesis framework.
Figure 1: Chain-of-Logic (highlighted part) organizes the rule application into a structured activity flow, enhancing the Domain-Specific Language (DSL) framework’s ability to handle complex tasks. The Neural Network Feedback Control mechanism (red path) utilizes data during synthesis to improve the performance of the synthesis process dynamically.
@@ 1 Introduction
In this paper, following the principles of fine-grained control and flexible modularity, we present COOL (Chain-Oriented Objective Logic), a neural-symbolic framework for complex program synthesis. At the core of our approach, we introduce the Chain-of-Logic (CoL), which integrates the functions of the activity diagram to enable fine-grained control Gomaa (2011). As illustrated in Figure 1, programmers can precisely organize rules into multiple activities and manage control flow using heuristics and keywords. Additionally, we leverage neural networks on top of CoL to dynamically fine-tune the synthesis process. For this purpose, we introduce Neural Network Feedback Control (NNFC) Turan & Jäschke (2024), which enhances future synthesis by learning from data generated during synthesis and suppresses neural network incorrect predictions through filtering. To ensure modularity, each neural network is bound with a specific CoL DSL, stored in separate library files for clear isolation and easy reuse. Thus, through the combination of CoL and NNFC, COOL achieves high efficiency and reliability when tackling complex synthesis tasks.
@@
Figure 3: Heuristic Vectors and Heuristic Values in Chain-of-Logic. Heuristic vectors h decompose the DSL $G$ into multiple sub-DSLs ($G_{0},G_{1},G_{2},G_{3}$) based on whether the value of the component at the corresponding position is non-zero. These sub-DSLs correspond to the activities depicted in Figure 1 and operate on partial programs $p$ using rules $r$. The synthesis process for each activity is guided by rule application policies $\pi_{r}$, which are generated by a heuristic algorithm $f$ that uses heuristic values $\textbf{h}[n]$ as input. In the experiments conducted in this paper, we adopt the A* algorithm and treat the heuristic value $\textbf{h}_{i}[n]$ as a reward for applying the rule $r_{i}$ during activity $n$ (Algorithm 1). Consequently, a higher heuristic value positively influences the rule’s application.
Figure 4: Keywords in Chain-of-Logic. In this illustrative CoL DSL, each node represents a stage or activity where a set of rules can be applied to generate partial programs. The flow between stages is managed by keywords return, logicjump(n), and abort, allowing for the implementation of complex control flow in program synthesis.
@@ 2.1 Chain-of-Logic (CoL)
Activity diagrams, widely used in software engineering, effectively describe how an initial state transitions to a final state through multiple activities. This feature aligns with the DSL-based program synthesis process. A DSL, defined as a context-free grammar, converts partial programs with nonterminal symbols into complete programs by applying given rules (Appendix A). However, as the rule set grows, DSL becomes inefficient in exploring partial programs. To enhance the efficiency of DSL, the Chain-of-Logic, drawing inspiration from activity diagrams, organizes the synthesis process into a sequence of manageable activities.
CoL improves the synthesis workflow of the DSL with two key features: heuristic vectors and keywords. As shown in Figure 3, heuristic vectors decompose DSL into multiple sub-DSLs that are consistent with the activity flow by dividing the rule applications scope. Within each activity, the DSL solver uses heuristic values to perform efficient program synthesis. For example, in Figure 1, a rule with the heuristic vector (0,7,3) belongs to sub-DSLs in activities 2 and 3 with heuristic values of 7 and 3, respectively. By dynamically pruning the search space and providing search guidance, heuristic vectors promote program synthesis efficiency.
Second, as illustrated in Figure 4, CoL introduces three keywords—return, logicjump(n), and abort—to dynamically manage state transition within or between activities during synthesis:
@@
    return: Ends the current rule, staying within current activity or advancing to following activities.
@@
    logicjump(n): Jumps directly to the activity $n$, enabling branching and loops within activity flow.
@@
Based on the principle of activity diagrams, CoL provides fine-grained control through heuristic vectors and keywords. This systematic approach enhances the efficiency of DSL synthesis (Refer to Appendix P for a detailed explanation of the CoL DSL synthesis process.).
Figure 5: Neural Network Feedback Control. The left side illustrates the complete control loop of NNFC. In the forward flow (green path), heuristic values $u$ guide the synthesis process as control signals. In the feedback loop (red path), the DSNN (Domain-Specific Neural Network, the neural network paired with a DSL) generates initial error signals $e_{0}$ from partial programs $y$. These singals are then filtered to produce high-quality error signals $e_{1}$, which adjust the initial heuristic values $u_{0}$. In multidomain synthesis, the CoL DSL and DSNN from the self-domain use partner domain information (dashed path) to clarify tasks and avoid competition, ensuring modularity. The right side details the feedback loop: The DSNN comprises multiple neural networks coupled in series via noise signals, with each network generating its own error signal $e_{0}$, then these signals with large discrepancies are filtered, retaining the final high-quality error signals $e_{1}$. (See Appendix B for details on the neural network’s format, architecture, training, prediction, and role in synthesis.)
@@ 2.2 Neural Network Feedback Control (NNFC)
The quality of the signals generated in the feedback loop directly determines the effectiveness of NNFC. If the error signals are of poor quality, NNFC may not only fail to provide additional improvements but also degrade CoL DSL performance. We ensure the error signal quality through an inner coupling structure within DSNN. As shown in Figure 5 (right), during synthesis tasks, DSNN processes partial programs using a series of sequentially connected neural networks. Each neural network takes both the partial programs and intermediate results from the preceding neural network as input, generating its own predictions. When errors occur in earlier networks, they propagate downstream as noise signals, amplifying at each stage. The difference in the outputs between these neural networks is positively correlated with the accumulated error. To mitigate this, we set a threshold to filter out signals with a significant difference in outputs. Finally, DSNN uses passed signals to generate multi-head outputs (Table 7) to fine-tune the forward flow:
@@ 3.1 Experimental Setup
Benchmarks. We evaluate CoL and NNFC using relational and symbolic tasks with varying difficulty levels, as outlined in Table 1. (Refer to Appendix C for specific task examples.) Specifically, the relational tasks are drawn from the CLUTRR Sinha et al. (2019) dataset, where the goal is to synthesize programs that capture specific target relationships based on human common-sense reasoning. In contrast, the symbolic tasks are generated by GPT Achiam et al. (2023). They involve synthesizing standard quadratic equation programs from non-standard quadratic forms by performing manual calculation steps. Although these tasks are simple for humans, they serve as a straightforward demonstration of how fine-grained control, derived from programmer expertise, can significantly improve program synthesis efficiency.
Metrics. Besides accuracy, we also focus on the following points: (1) CPU Overhead is assessed by the number of tree operations required for synthesis. (2) Memory overhead is assessed by the number of transformation pairs (a partial program paired with the rule to be applied. Appendix A.3).(3) GPU Overhead is measured by the number of neural network invocations. (4) Time overhead is referenced by the actual time spent on program synthesis tasks. (5) Filtering Performance is evaluated by the attenuation ratio of invalid to passed neural network predictions.  [^1] Resource limit: Each partial program must be completed with a cost of at most 1000 transformation pairs and the length of the synthesis path cannot exceed 50.
@@
Table 2: CoL DSL configurations. The DSL for relational benchmarks has a limited search space and shorter CoL, facing challenges from numerous production rules leading to larger trees. Conversely, the DSL for symbolic benchmarks offers an unlimited search space with a longer CoL, but the many permutation rules increase the risk of cyclic rule applications. (Refer to Appendices I and J for source code snippets of the CoL DSLs used in experiments.)
@@ 3.3 Dynamic Experiments
Further ablation experiments confirm that reliability provided by NNFC primarily stems from the filtering effect of the inner coupling structure. As shown in Figures 7 and 9, the inner coupling structure reduces the occurrence of accuracy declines due to DSNN mispredictions by 94%. Additionally, the dynamic performance reveals how the inner coupling structure enhances NNFC:
In the scenarios where a DSNN underperforms due to issues such as insufficient training data Mikołajczyk & Grochowski (2018) (as seen in Figure 7, tasks 51-100), inadequate generalization to more challenging tasks Yosinski et al. (2014); Wei et al. (2019) (Figure 7, tasks 301-350), and catastrophic forgetting when tasks from a new domain are learned Kirkpatrick et al. (2017); Van de Ven & Tolias (2019) (Figure 9, tasks 1-100), incorrect predictions lead the actual synthesis path to deviate from the CoL, which in turn causes inefficiency and reduced accuracy. During these phases, for NNFC with the inner coupling structure, the attenuation ratio spikes, indicating that a large percentage of neural network predictions are filtered out. Consequently, the inner coupling structure ensures that the synthesis process adheres to the CoL, effectively mitigating the negative impact of DSNN mispredictions and enhancing reliability.
@@ 4 Conclusion
Appendix A DSL Program Synthesis in COOL
COOL adopts a top-down synthesis strategy that converts input partial programs with nonterminals into complete programs by applying a sequence of well-defined transformation rules.
A.1 Input Program
In the relational reasoning tasks, the input is COOL code such as Code A.1:
Code A.1: Relational Reasoning Task Input Program (Wesley) is (James)s son & (Martha) is (Wesley)s daughter & (Hugh) is (Martha)s uncle & (Hugh) is (James)s (\$relation);
where \$ specifies the nonterminal, indicating that the DSL solver needs to synthesize a complete program that calculates the correct value for relation (the relationship between Hugh and James) in order to satisfy the given specification.
The symbolic task input is The input for symbolic tasks is as follows (Code A.1):
Code A.2: Symbolic Reasoning Task Input Program \$x^2 + 4*\$x == 3;
Similarly, the DSL solver needs to generate a complete program that calculates the value of x.
A.2 Output Program
As shown in Code A.2, for family relationship reasoning tasks, the synthesized program is:
Code A.3: Relational Reasoning Task Output Program relation = "son";
For symbolic reasoning tasks, the generated output program is shown in Code :
Code A.4: Symbolic Reasoning Task Output Program Invoke Quadratic Solution Formula( a=1, b=4, c=-3, x ); // a, b, c are coefficients in "1*x^2 + 4*x + (-3) == 0"
In reality, the program synthesis takes place at the intermediate representation level (see Appendix P), and Codes A.2 A.2 are provided for explanatory purposes.
A.3 Traditional DSL
A DSL, defined as a context-free grammar:
$G=\{V,\Sigma,R,S\},$ (1)
where $V$ is the set of non-terminal symbols, $\Sigma$ is the set of terminal symbols, $R$ is the set of rules, and $S$ is the starting symbol (in this context, it is a partial program). The DSL’s derivation process converts partial programs with nonterminal symbols into complete programs by applying given rules.
The synthesis process for traditional DSLs involves iteratively transforming partial programs into complete programs by applying a series of rules. Each partial program ($p$) and the corresponding rule ($r$) to be applied to it form a transformation pair $(p,r)$. When a rule is applied, it modifies the syntax tree of the partial program through what we refer to as a tree operation. The synthesis process consists of a series of transformation pairs connected by tree operations, which is known as a synthesis path or trajectory. These paths are classified into three kinds:
  • Feasible Path: Leads to a complete program.
  • Infeasible Path: Proven to be unable to synthesize a complete program.
  • Unfinished Path: Still in progress.
To clarify key concepts involved in the synthesis process, we provide the following definitions of terms:
  • Tree Operation/Manipulation: Refers to the modification of the syntax tree of a partial program during the synthesis process. It is essential for transforming partial programs into complete programs and has an associated CPU cost.
  • Transformation Pair $(p,r)$: A combination of a partial program and a rule to be applied. It records the explored space and possible exploration directions, requiring memory storage.
  • Synthesis Path/Trajectory: A sequence of transformation pairs, $\{(p_{0},r_{0}),(p_{1},r_{1}),\dots\}$, representing the process of transforming a partial program into a complete one. Its function is to track the entire synthesis process, whether it leads to a feasible, infeasible, or unfinished path.
A.4 CoL DSL
Compared with traditional DSLs that apply rules to input programs without a clear destination to synthesize output programs, the Chain-of-Logic (CoL) allows the programmer to outline the flow of activities to synthesize the complete program from an initial partial program. For example, in Figure 3, the activity flow is:
Start $\rightarrow$ 1 Separate Relations and Genders $\rightarrow$ 2 Reason Inverse Relations $\rightarrow$ 3 Reason Indirect Relations $\rightarrow$ 4 Recombine Relations and Genders, Eliminate Irrelevant Relations $\rightarrow$ End.
Each activity in the synthesis process has a corresponding sub-DSL decomposed from the original DSL for transforming the program from one state to another.
Therefore, a CoL DSL with $n$ activities can be defined as multiple sub-DSLs in series:
CoL $\text{CoL }G=\{G_{1},G_{2},\dots,G_{n}\}$ (2)
A sub-DSL for activity $i$ is defined as:
$G_{i}=\{V,\Sigma,\{(r,\textbf{h},k)\mid\textbf{h}[n]\neq 0\text{ and }r\in R\},S,f\}$ (3)
where $\textbf{h}=(h_{1},h_{2},\dots)$ in $(r,\textbf{h},k)$ represents the heuristic vector bound with $r$, and the components are termed as heuristic values. $h[n]$ represents the $n$-th component of $h$, which is the effective heuristic value of rule $r$ in activity $n$. $\textbf{h}[n]$ is a parameter of the sub-DSL’s program synthesis algorithm $f$. It guides the direction of program synthesis by affecting the application decisions of the rules bound to it, thereby improving synthesis efficiency and accuracy. The specific role of the heuristic value is determined by the search algorithm $f$ used by the DSL program synthesis. To conduct controlled variable experiments, we regard the heuristic value as a reward (negative cost) and use the A* algorithm as the search algorithm for all DSLs, which means that in activity $n$, under the same circumstances, the rule with a larger $\textbf{h}[n]$ will be applied first, and the derived program will also be considered promising and will be explored further with priority.
$k$ represents keyword(s) in a rule’s specific logic, which controls the program state transition within an activity or between activities.
The Chain-of-Logic provides a comprehensive methodology for achieving fine-grained control over DSL program synthesis. This approach allows programmers to explicitly break down the synthesis process into distinct phases or activities, with each activity corresponding to a specific sub-DSL.
A.5 A* Search in Program Synthesis